OpenStudy (anonymous):

line ab represented by the equation y=3x-2 line cd is parallel to line ab and passes through the point (5,2) what is the equation of lined cd in slope intercept form

OpenStudy (anonymous):

y=mx+c insert (5,2) and m=3 (as it is parallel) 2=3(2)+c c=-4 answer y=3x-4 (is in intercept form)

OpenStudy (anonymous):

my mistake y=mx+c insert (5,2) and m=3 (as it is parallel) 2=3(5)+c c=-13 answer y=3x-13 (is in intercept form)